# OMR-Diffusion:Optimizing Multi-Round Enhanced Training in Diffusion Models for Improved Intent Understanding

### v1

@@ Abstract
Generative AI has significantly advanced text-driven image generation, but it still faces challenges in producing outputs that consistently align with evolving user preferences and intents, particularly in multi-turn dialogue scenarios. In this research, We present a Visual Co-Adaptation (VCA) framework that incorporates human-in-the-loop feedback, utilizing a well-trained reward model specifically designed to closely align with human preferences. Using a diverse multi-turn dialogue dataset, the framework applies multiple reward functions (such as diversity, consistency, and preference feedback) to refine the diffusion model through LoRA, effectively optimizing image generation based on user input. We also constructed multi-round dialogue datasets with prompts and image pairs that well-fit user intent. Experiments show the model achieves 508 wins in human evaluation, outperforming DALL-E 3 (463 wins) and others. It also achieves 3.4 rounds in dialogue efficiency (vs. 13.7 for DALL-E 3) and excels in metrics like LPIPS (0.15) and BLIP (0.59). Various experiments demonstrate the effectiveness of the proposed method over state-of-the-art baselines, with significant improvements in image consistency and alignment with user intent. The project page is https://tathataai.github.io/OMR-Diffusion/.
@@ 1 Introduction
A wide range of industries benefit from the application of generative artificial intelligence, as it allows for the optimization of both creative and non-creative tasks. Several innovative models, including DALL·E 2 [1], Imagen [2], and Stable Diffusion [3] have shown extraordinary capabilities in generating unique, convincing, and lifelike images from textual descriptions. Through these models, users are able to visualize ideas that were previously difficult to materialize, opening up new opportunities in content creation, design, advertising, and entertainment [4, 5, 6, 7]. Despite this, these models still face significant challenges in meeting the diverse needs of users. An important limitation is the interpretation of differentiated user preferences, particularly capturing changes in personal style and intentions over time [8, 9]. Many existing models operate statically, producing images based on fixed prompts without incorporating user feedback into subsequent generations [10]. This disconnects user expectations from generated outputs, particularly for applications requiring iterative refinement or personalized results. Furthermore, current generative models have difficulty maintaining consistency across rounds of multi-turn interactions, which can result in outputs that deviate from the user’s intentions during the conversation [11, 12]. Consequently, non-experts find it difficult to fine-tune results due to the opacity of variable adjustments  [13, 14]. Finally, while some frameworks integrate reinforcement learning (RL) to align model outputs with human preferences, these approaches are not always effective in capturing the subtleties of individual user feedback. Especially for creative tasks that require high semantic fidelity and personalization, using reinforcement learning alone may lead to suboptimal alignment [15]. These challenges highlight the urgent need for more dynamic, feedback-driven, and generative AI systems that can iteratively adapt to user preferences and simplify interactions between experts and non-experts. [16, 17].
Figure 1: The workflow demonstrates how human preferences guide text-to-image diffusion, with a DPO-trained reward model evaluating image-prompt alignment and PPO updating LoRA parameters while keeping the diffusion model fixed.
To tackle these challenges, we propose an innovative approach that enhances the user experience by simplifying the interaction between users and generative models. As opposed to traditional models that require extensive knowledge of the control elements, our approach is designed to take advantage of the concept of "human in the loop" adaptation [10, 18]. By integrating continuous user feedback, our model dynamically evolves to better meet user expectations. This feedback-driven adaptation not only improves the accuracy and relevance of the generated images but also reduces the learning curve for users, making the technology more accessible and empowering. Figure 1 depicts the overall workflow of our model, demonstrating how it incorporates user feedback to iteratively refine image generation. Our key contributions are summarized below:
@@ 2 Related Work
In recent years, artificial intelligence technology has achieved breakthrough development in multiple disciplines, covering multiple dimensions such as low-level visual processing  [7, 19, 20, 21, 22], high-level visual applications[23, 24, 25], intelligent engineering systems [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and image generation technology [39, 40, 41, 42].
@@ 2.1 Text-Driven Image Editing Framework
Image editing through textual prompts has revolutionized how users interact with images, making editing processes more intuitive. One of the seminal works is Prompt-to-Prompt (P2P) [43]. The core idea behind P2P is aligning new information from the prompt with the cross-attention mechanism used during the image generation process. P2P allows modifications to be made without retraining or adjusting the entire model, but simply by changing attention maps in specific areas. Expanding upon P2P, MUSE [44] introduced a system that allows both textual and visual feedback during the generation process. The enhancement made the system more adaptive, enabling it to better respond to user inputs, regardless of whether the input was provided as text or through visual corrections. Building on these developments, the Dialogue Generative Models framework [11] took the concept of text-driven editing further by integrating dialogue-based interactions. It allows for a more fluid conversation between the user and the model, iteratively refining the generated image through feedback loops, improving alignment with user preferences over multiple interactions. In another advancement, Prompt Charm [17] introduced techniques that refine prompt embeddings to provide more precise control over specific image areas without necessitating retraining of the entire model. More recently, ImageReward [13] has adopted a different approach by utilizing human feedback to optimize reward models. Iterative adjustments are made to image outputs to align them more closely with human preferences, emphasizing the improvement of text-to-image coherence [45, 46].
@@ 4.3 Comparison Study
This Figure  5 compares several generative models, including SD v1.4, P2P, Muse, DALL-E 3, Imagen, and Cogview 2, using two key evaluation metrics - user preference scores (top panel) and CLIP scores (bottom panel). User preference scores reflect subjective human evaluations with wider interquartile ranges suggesting greater discernability between models. Considering such variability, preference scores may effectively capture subtle variations in user satisfaction and alignment with human judgments. CLIP scores, which measure semantic alignment between text and generated images, show narrower interquartile ranges and medians, suggesting limited sensitivity to subtle changes in generative quality. It is noteworthy that models such as Muse and Cogview 2 perform well when measured against these metrics, while user preference scores highlight the strengths of other models, such as Imagen and DALL-E 3. This disparity underscores the importance of incorporating human evaluations alongside automated metrics like CLIP, as the latter may fail to fully account for subjective aspects such as aesthetic appeal and intent alignment.
@@
This Figure  6 compares the win, tie, and lose rates of our model against four different selection methods: Random selection (randomly selecting 3 images), CLIP Score, Aesthetic Score, and BLIP Score [64], across three scenarios where the top 3 images are selected from 9, 25, and 64 images, respectively. Across all scenarios, our model’s preference score consistently outperforms the other methods with win rates exceeding 50%. In addition, the performance of our model remains stable with increasing numbers of images, indicating the robustness of our approach in capturing human intent effectively. As compared to CLIP Score, Aesthetic Score, and BLIP Score, our preference score has a higher win rate, demonstrating its superior alignment with human judgments. Hence, our approach generates selections that better reflect the preferences of users.
@@
This heatmap Figure  7 compares the win rates of various generative models, including SD v1.4, DALL·E 3, P2P, Imagen, Muse, CogView 2, and Ours, across 8 rounds of dialogue interactions. Each cell represents the win rate of the model in the row compared to the model in the column, with higher values (closer to red) indicating better performance.
@@ 4.4 Ablation Study
This Figure  8 demonstrates the impact of removing consistency, user alignment, and diversity rewards on generating images from the prompt: "Modify this figure to a cartoon style with multiple apples and a happy squirrel." Without the diversity reward, images become repetitive and less varied, resulting in a lack of creativity. Without the consistency reward, the outputs become incoherent, deviating from the intended style and composition, as seen in the replacement of the squirrel with unrelated elements. Excluding the mutual information reward causes the generated image to lose its connection to the prompt, resulting in outputs that fail to capture the desired theme or maintain semantic alignment. In the final output produced by our model (rightmost column), all rewards have been successfully incorporated, resulting in an accurate, coherent, and stylistically consistent image that matches the user’s instructions.

### v3

@@ Abstract
Deep learning has made impressive progress in natural language processing (NLP), time series analysis, computer vision, and other aspects [1, 2, 3, 4, 5, 6]. Generative AI has significantly advanced text-driven image generation, but it still faces challenges in producing outputs that consistently align with evolving user preferences and intents, particularly in multi-turn dialogue scenarios. In this research, We present a Visual Co-Adaptation (VCA) framework that incorporates human-in-the-loop feedback, utilizing a well-trained reward model specifically designed to closely align with human preferences. Using a diverse multi-turn dialogue dataset, the framework applies multiple reward functions (such as diversity, consistency, and preference feedback) to refine the diffusion model through LoRA, effectively optimizing image generation based on user input. We also constructed multi-round dialogue datasets with prompts and image pairs that well-fit user intent. Experiments show the model achieves 508 wins in human evaluation, outperforming DALL-E 3 (463 wins) and others. It also achieves 3.4 rounds in dialogue efficiency (vs. 13.7 for DALL-E 3) and excels in metrics like LPIPS (0.15) and BLIP (0.59). Various experiments demonstrate the effectiveness of the proposed method over state-of-the-art baselines, with significant improvements in image consistency and alignment with user intent. The project page is https://tathataai.github.io/OMR-Diffusion/.
@@ 1 Introduction
Generative artificial intelligence (AI) has demonstrated transformative potential across industries by optimizing both creative workflows and routine tasks. Pioneering models like DALL·E 2 [7] and Stable Diffusion [8] have revolutionized text-to-image conversion, yet persistent challenges in semantic alignment become evident when handling text-rich scenarios. Recent benchmarks like [9, 10] reveal that 68% of text-visual inconsistencies originate from inadequate layout-text synchronization - a limitation also observed in document understanding systems. The work of [11] demonstrates how tokenized layout representations can improve spatial-textual alignment by 23%, while [12]’s frequency domain analysis achieves 89.7% accuracy in OCR-free document parsing through multimodal fusion.
Three critical limitations hinder broader adoption:
Dynamic Intent Capture: Current models struggle with evolving user preferences, particularly in multilingual contexts. [13]’s ego-evolving recognizer adapts through in-context learning, yet [14]’s evaluation across 12 languages still shows 41
Conceptual Coherence: While [15]’s harmonized architecture improves text-visual synchronization by 32
Interactive Refinement: The trial-and-error process identified in [16] persists, where [17]’s RL-based box adjustment reduces annotation iterations by 58%, though [18]’s single-point spotting demonstrates 76.2% accuracy with weak supervision.
Emerging solutions show promise but face new challenges. [19]’s feature sampling strategy proves sparse features outperform dense computation in complex text detection (89.7% F-score), while [20]’s partial-global view fusion enhances layout control by 19%. Reinforcement learning approaches, while effective for basic alignment [21], achieve only 67% satisfaction in creative tasks requiring high semantic fidelity. [22]’s visual instruction tuning framework improves multi-turn consistency by 32%, yet struggles with complex layout control - a limitation addressed partially by [23, 24]’s audio annotation system. The integration of these advances suggests a path forward: hybrid systems combining [11]’s layout awareness with [25, 25]’s concept synergy could bridge the remaining 23% performance gap in text-rich generation tasks.
@@ 2 Related Work
In recent years, artificial intelligence technology has achieved breakthrough development in multiple disciplines, covering multiple dimensions such as low-level visual processing  [26, 27, 28], high-level visual applications[29, 30, 31], intelligent engineering systems [32, 33, 34, 35, 36, 37, 38, 39, 40], and image generation technology [41, 42, 43].
@@ 2.1 Text-Driven Image Editing Framework
Image editing through textual prompts has revolutionized how users interact with images, making editing processes more intuitive. One of the seminal works is Prompt-to-Prompt (P2P) [44]. The core idea behind P2P is aligning new information from the prompt with the cross-attention mechanism used during the image generation process. P2P allows modifications to be made without retraining or adjusting the entire model, but simply by changing attention maps in specific areas. Expanding upon P2P, MUSE [45] introduced a system that allows both textual and visual feedback during the generation process. The enhancement made the system more adaptive, enabling it to better respond to user inputs, regardless of whether the input was provided as text or through visual corrections. Building on these developments, the Dialogue Generative Models framework [46] took the concept of text-driven editing further by integrating dialogue-based interactions. It allows for a more fluid conversation between the user and the model, iteratively refining the generated image through feedback loops, improving alignment with user preferences over multiple interactions. In another advancement, Prompt Charm [47] introduced techniques that refine prompt embeddings to provide more precise control over specific image areas without necessitating retraining of the entire model. More recently, ImageReward [48] has adopted a different approach by utilizing human feedback to optimize reward models. Iterative adjustments are made to image outputs to align them more closely with human preferences, emphasizing the improvement of text-to-image coherence [49].
@@ 4.3 Comparison Study
This Figure  4 compares several generative models, including SD v1.4, P2P, Muse, DALL-E 3, Imagen, and Cogview 2, using two key evaluation metrics - user preference scores (top panel) and CLIP scores (bottom panel). User preference scores reflect subjective human evaluations with wider interquartile ranges suggesting greater discernability between models. Considering such variability, preference scores may effectively capture subtle variations in user satisfaction and alignment with human judgments. CLIP scores, which measure semantic alignment between text and generated images, show narrower interquartile ranges and medians, suggesting limited sensitivity to subtle changes in generative quality. It is noteworthy that models such as Muse and Cogview 2 perform well when measured against these metrics, while user preference scores highlight the strengths of other models, such as Imagen and DALL-E 3. This disparity underscores the importance of incorporating human evaluations alongside automated metrics like CLIP, as the latter may fail to fully account for subjective aspects such as aesthetic appeal and intent alignment.
@@
This Figure  5 compares the win, tie, and lose rates of our model against four different selection methods: Random selection (randomly selecting 3 images), CLIP Score, Aesthetic Score, and BLIP Score [69], across three scenarios where the top 3 images are selected from 9, 25, and 64 images, respectively. Across all scenarios, our model’s preference score consistently outperforms the other methods with win rates exceeding 50%. In addition, the performance of our model remains stable with increasing numbers of images, indicating the robustness of our approach in capturing human intent effectively. As compared to CLIP Score, Aesthetic Score, and BLIP Score, our preference score has a higher win rate, demonstrating its superior alignment with human judgments. Hence, our approach generates selections that better reflect the preferences of users.
@@
This heatmap Figure  6 compares the win rates of various generative models, including SD v1.4, DALL·E 3, P2P, Imagen, Muse, CogView 2, and Ours, across 8 rounds of dialogue interactions. Each cell represents the win rate of the model in the row compared to the model in the column, with higher values (closer to red) indicating better performance.
@@ 4.4 Ablation Study
This Figure  7 demonstrates the impact of removing consistency, user alignment, and diversity rewards on generating images from the prompt: "Modify this figure to a cartoon style with multiple apples and a happy squirrel." Without the diversity reward, images become repetitive and less varied, resulting in a lack of creativity. Without the consistency reward, the outputs become incoherent, deviating from the intended style and composition, as seen in the replacement of the squirrel with unrelated elements. Excluding the mutual information reward causes the generated image to lose its connection to the prompt, resulting in outputs that fail to capture the desired theme or maintain semantic alignment. In the final output produced by our model (rightmost column), all rewards have been successfully incorporated, resulting in an accurate, coherent, and stylistically consistent image that matches the user’s instructions.